How many liters of a 20% acid solution must be mixed with a 60% solution to obtain 40 liters of a 35% solution?
Answer by stanbon(75887)

How many liters of a 20% acid solution must be mixed with a 60% solution to obtain 40 liters of a 35%
---------------------
Equation:
acid + acid = acid
0.20x + 0.60(40-x) = 0.35*40
----
Multiply thru by 100 to get:
20x + 60*40 - 60x = 35*40
-40x = -25*40
x = 25 liters (amt. of 20% solution needed in the mixture)
----
Cheers,
Stan H.